Justify whether it is possible to design an expanding opcode to allow the following to be encoded in 16-bit instruction? An address is 4 bits.


15 instructions with three addresses

14 instructions with two addresses

31 instructions with one address

18 instructions with zero address


1
Expert's answer
2021-07-05T04:25:58-0400

For the first 15 instructions with 3 addresses

15×2^4×2^2×4= 61440 bit patterns

For the second 14 instruction with 2 addresses

14×2^4×2^4= 3584 bit patterns

For the third 31 instruction with address

31×2^4= 496 bit patterns

For the 18 instruction represents 18 bit patterns

Required bit pattern is 61440+3584+496+18= 65538

With 16 bit we can only have

2^16= 65536

Required bit pattern (65 538)is more than what we have (65 536) so it is not possible.